# Gibbs sampling, exact sampling, and distribution function evaluation for the generalized inverse Gaussian distribution

### v2

@@ Appendix A Proofs of Propositions
$\mathcal{Z}_{0}=\frac{b^{(p+1)/2}2^{1/2}\Gamma(-p-1/2)K_{p}(a^{1/2}b^{1/2})}{a^{p/2}\pi^{1/2}}.$
If $p>-1/2$:
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{X}(x\mid y)}{f_{Y}(y\mid x)}\,\mathrm{d}x$
$=\int_{0}^{\infty}\frac{\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y)}{\mathrm{Gamma}(t\mid p+1/2,1/x)}\,\mathrm{d}x$
where
Proposition 2.
Proof.
$f_{X,Y}(x,y)\propto\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,y^{-(p+3/2)}\,\exp\left(-xy\right),$
from which it is straightforward to deduce that
$Y\mid X=x$ $\sim\mathrm{Gamma}(-(p+1/2),x).$
If $p>-1/2$, then
$f_{X,Y}(x,y)\propto y^{p-1/2}\exp\left\{-\sqrt{a(b+2y)}\right\}\,\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y),$
which implies that
$Y\mid X=x$ $\sim\mathrm{Gamma}(p+1/2,1/x).$
Proposition 3.
Proof.
As we stated in the main text, we prove this result by checking that the conditions of Theorem 3.5 in [21] are satisfied. The proof amounts to showing that there exist functions $f,g:[0,\infty)\rightarrow[1,\infty)$ and constants $j,k,m,n>0$ with $jm<1$ so that
Case $p<-1/2$: Define
$\theta$ $=3/(2b)$
$\mu$ $=3(1/2-p)/(2b)$
$f(x)$ $=1+\alpha(1/x-\theta)^{2},\qquad 0<\alpha<b^{2}/(p^{2}-1/4)$
$\mathrm{Var}(1/X\mid Y=y)$ $={(a+2y)^{1/2}}{b^{-3/2}}+{2}{b^{-2}}$
Rearranging terms, we obtain
$=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b+2\alpha(y-\mu)/b$
$\leq 1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b+g(y),$
$\mathbbm{E}[f(X)\mid Y=y]$ $\leq jg(y)+k$
$j$ $=1$
$k$ $=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b.$
Both $j$ and $k$ are positive, which is required by Theorem 3.5 in [21]. Now, we find the conditional expectation of $g$ and bound it:
$\leq 1+9\beta(1/2-p)/(4b^{2})+\beta(p^{2}-1/4)f(x)/\alpha.$
We can rewrite the inequality as
$\mathbbm{E}[g(Y)\mid X=x]$ $\leq mf(x)+n$
$m$ $=\beta(p^{2}-1/4)/\alpha$
$n$ $=1+9\beta(1/2-p)/(4b^{2}),$
where $m,n>0$, as required by Theorem 3.5 in [21]. It remains to show that $jm<1$. Given our choices of $\alpha$ and $\beta$,
$10<{\beta(p^{2}-1/4)}/{\alpha}<$ 1
Case $p>-1/2:$ Let $10<\gamma<$ 1 and
$\theta$ $=1/(2a)$
$\mu$ $=(p+1/2)/(2a)$
$f(x)$ $=1+\alpha(x-\theta)^{2},\qquad\alpha>0$
$=1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a+{2\alpha}(y-\mu)/a$
$\leq 1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a+\gamma g(y).$
$\mathbbm{E}[f(X)\mid Y=y]$ $\leq jg(y)+k$
$j$ $=\gamma$
$k$ $=1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a,$
where $j$ and $k$ are positive, as required by Theorem 3.5 in [21]. Finally, we bound the conditional expectation of $g$:
$=1+\beta\left\{(p+1/2)x^{2}+[(p+1/2)x-\mu]^{2}\right\}$
$=1+\beta(p+1/2)x^{2}+\beta(p+1/2)^{2}(x-\theta)^{2}$
$=\beta(p+1/2)x^{2}+[\beta(p+1/2)^{2}-\alpha](x-\theta)^{2}+f(x)$
$=Q(x)+f(x)$
$\leq c_{0}+f(x),$
$\mathbbm{E}[g(Y)\mid X=x]$ $\leq mf(x)+n$
$m$ $=1$
$n$ $=c_{0},$
Proposition 4.
$f_{Y}(y)$ $=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b)$
$w$ $={K_{p-2}(\sqrt{ab})}/{K_{p}(\sqrt{ab})}\in(0,1)$
and $E\sim\mathrm{Exp}(a/2).$
Proof.
$f_{X}(x)$ $=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y$
where
$f_{Y}(y)$ $=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b)$
$w$ $=\frac{K_{p-2}(\sqrt{ab})}{K_{p}(\sqrt{ab})}.$
Let $a>0,b>0,p>1$, and define:
[…]
The ratio
[…]
is monotone non-decreasing in $x>0$ for $p>1$, so we can apply Theorem 2.4 in [20] to write $g(x)$ as a mixture of truncated $\mathrm{Exp}(a/2)$ random variables:
$g(x)=\mathrm{GIG}(x\mid p,a,b)=\int_{0}^{\infty}\frac{a}{2}\exp\left\{-\frac{a}{2}(x-y)\right\}\mathbbm{1}(x\geq y)\,f_{Y}(y)\,\mathrm{d}y.$
By Theorem 2.4 in [20], the mixing density $f_{Y}(y)$ is
[…]
$=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b).$
$f_{\tilde{X}}(x)$ $=\int_{0}^{\infty}f_{E}(x-y)f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{\infty}\mathrm{Exp}(x-y\mid a/2)\,f_{Y}(y)\,\mathrm{d}y$
$=\mathrm{GIG}(x\mid p,a,b),$
as we wanted to show. ∎
Proposition 5.
where
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function [1].
Proof.
Let $G_{p}$ be the CDF of $\mathrm{GIG}(p,a,b)$. Then,
On the one hand,
$\int_{0}^{x}f_{Y}(y)\,\mathrm{d}y=wG_{p-2}(x)+(1-w)G_{p-1}(x),$
$\int_{0}^{x}\exp\left(\frac{ay}{2}\right)f_{Y}(y)\,\mathrm{d}y=I_{p}(x)$
where $\Gamma(\cdot,\cdot)$ is the incomplete gamma function [1]. ∎

### v3

@@ B.2 Propositions in the new parametrization
$I_{p}(x)=\frac{\tau\,\omega^{\,p-2}}{2^{\,p-1}K_{p-2}(\omega)}\Gamma\left(2-p,\frac{\omega\eta}{2x}\right)\;+\;\frac{(1-\tau)\,\omega^{\,p-1}}{2^{\,p}K_{p-1}(\omega)}\Gamma\left(1-p,\frac{\omega\eta}{2x}\right),$
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function.
B.3 Gibbs sampler based on Zhang and Reiter (2022)
This section describes an alternative Gibbs sampler for GIG that uses a decomposition presented in Zhang and Reiter (2022). The Gibbs sampler iterates between sampling from truncated distributions and the support of the conditionals changes from iteration to iteration. This characteristic can introduce challenges in verifying certain convergence properties, as discussed later.
B.3.1 Notation for Truncated Distributions
We denote truncated distributions using the following notation:
  • $X\sim\text{TInvGamma}(\alpha,\beta,L,U)$: a truncated inverse gamma distribution with shape parameter $\alpha$, scale parameter $\beta$, and truncation bounds $(L,U)$.
  • $Y\sim\text{TExp}(\lambda,L,\infty)$: a truncated exponential distribution with rate $\lambda$ and lower truncation bound $L$.
Truncated densities are normalized over their truncation bounds.
B.3.2 Joint Distribution
Using results in Zhang et al. (2012), we can write a joint distribution of $X$ and $Y$ for $X\sim\text{GIG}(p,a,b)$:
$f_{X,Y}(x,y)\propto\exp\left(-\frac{by}{2}\right)\,x^{p-1}\exp\left(-\frac{a}{2x}\right),\quad 0<x<y,$
where $p\leq 0$.
From this joint distribution, the two full conditional distributions are derived by isolating terms involving $x$ and $y$, respectively.
B.3.3 Full Conditional Distributions
From $f_{X,Y}(x,y)$, the terms involving $x$ give:
$f_{X\mid Y}(x\mid y)\propto x^{p-1}\exp\left(-\frac{a}{2x}\right),\quad 0<x<y.$
This corresponds to:
$X\mid Y=y\sim\text{TInvGamma}(-p,a/2,0,y).$
Similarly, the terms involving $y$ yield:
$f_{Y\mid X}(y\mid x)\propto\exp\left(-\frac{by}{2}\right),\quad y>x.$
This corresponds to:
$Y\mid X=x\sim\text{TExp}(b/2,x,\infty).$
B.3.4 Generalization for all $p$
For $p>0$, we can use the property $1/X\sim\text{GIG}(-p,b,a)$ to extend the sampler:
  1. 1.
    Transform $X$ to $Z=1/X$ and apply the Gibbs sampler for $Z\sim\text{GIG}(-p,b,a)$.
  2. 2.
    After sampling $Z$, invert to obtain $X=1/Z$.
B.3.5 Assessing convergence
Theorem 1 in Johnson and Burbank (2015) provides sufficient conditions for geometric ergodicity of two-stage Gibbs samplers. In this case, the support of the transition kernel changes at every iteration, which complicates the verification of Assumption $\mathcal{A}$ in Johnson and Burbank (2015). It is known that changing the support of the transition kernel from iteration to iteration can negatively affect the properties of the sampler (Robert and Casella, 2004; Meyn and Tweedie, 2012).
B.4 Numerical experiments
We conducted two experiments to compare the efficiency of the data-augmented Gibbs sampler and the exact simulation method with the methods proposed by Devroye (2014), as implemented in the R package boodist (Laurent, 2023), and Hörmann and Leydold (2014), as implemented in the R package GIGrvg (Leydold and Hormann, 2023). We compared with the methods from these articles because, as far as we are aware, they represent the two most recent published works on simulating from the generalized inverse Gaussian distribution. [In the more recent preprint of Zhang and Reiter (2022), the authors mention that their method never outperformed that of Hörmann and Leydold (2014).] Before we describe the experiments, it is important to note that the efficiency of a method is highly dependent upon the details of its implementation, including the programming language, the skill of the programmer, etc.
In the first experiment, we compared the two Gibbs samplers described in the example toward the end of Section 2. The original Gibbs sampler alternates between normal and GIG full conditionals, while the data-augmented Gibbs sampler cycles through normal, inverse Gaussian, and gamma full conditionals. We considered two versions of the original Gibbs sampler: one using the method of Devroye (2014) implemented in boodist and one using the method of Hörmann and Leydold (2014) implemented in GIGrvg. The data were simulated from the normal distribution with parameters $\mu=1,\sigma^{2}=1$ and sample size $n=100.$ The prior hyperparameters were chosen as $\theta_{0}=0,\tau_{0}^{2}=100,p_{0}=3/4,a_{0}=1,b_{0}=1.$ We ran the Gibbs samplers for 5000 iterations. Based on 1000 repetitions, we found that the data-augmented Gibbs sampler was about 8% slower than the version of the original Gibbs sampler using the method of Hörmann and Leydold (2014) implemented in GIGrvg. However, the data-augmented Gibbs sampler was more than 9 times faster than the version of the original Gibbs sampler using the method of Devroye (2014) implemented in boodist. Again, these differences may tell us more about the quality of the implementations of each method than they tell us about the methods themselves.
Neither the original Gibbs sampler nor the data-augmented Gibbs sampler produce independent draws from the posterior distribution, so it is important to evaluate the effective sample size (ESS) in addition to the time it takes to produce the Markov chains. For the original Gibbs sampler, we found that the ESS was close to the nominal sample size for both parameters $\mu$ and $\sigma^{2}.$ For the data-augmented Gibbs sampler, this was only true for $\mu.$ The ESS of $\sigma^{2}$ was about 1/3 of the nominal sample size. To assess the ESS, we simulated 100 chains of length 50000 with 5000 warmup iterations and used the effectiveSize function from the R package coda (Plummer et al., 2006).
The conclusions of the previous two paragraphs did not seem to be sensitive to the particular choices of parameters, sample size, or prior hyperparameters. A summary of the results of the first simulation experiment appears in Table 1.
Table 1: A summary of the results of the first simulation experiment. The “Time” column provides the cumulative time in seconds to produce 1000 chains of length 5000. The “Relative” column is computed by dividing the “Time” column by its smallest value. The last two columns report the median ESS per iteration of $\mu$ and $\sigma^{2}$ from 100 chains of length 50000 with 5000 warmup iterations. We do not report ESS per iteration for the method of Devroye (2014) because it should not differ from that of the method of Hörmann and Leydold (2014).
Sampler Time Relative ESS/iter $\mu$ ESS/iter $\sigma^{2}$
Proposed 26.564 1.085 1.00 .339
L & H 24.491 1.000 1.00 .981
Devroye 228.304 9.322 * *
In the second experiment, we compared the exact simulation method of Section 3 with the methods of Devroye (2014) and Hörmann and Leydold (2014) in the case where $p=3/2.$ The purpose of the experiment was to establish that the exact simulation method outperforms the other methods when $|p|$ is sufficiently small. We set $p=3/2,a=1,b=1.$ Based on $10000$ repetitions, we found that the method of Hörmann and Leydold (2014) implemented in GIGrvg took almost 10% longer than the exact simulation method of Section 3. The method of Devroye (2014) implemented in boodist took about 87% longer than the exact simulation method. These conclusions did not seem to be sensitive to the particular choices of the parameters $a$ and $b.$ A summary of the results of the second simulation experiment appears in Table 2.
Table 2: A summary of the results of the second simulation experiment. The “Time” column provides the cumulative time in seconds to complete all 10000 repetitions. The “Relative” column is computed by dividing the “Time” column by its smallest value.
Sampler Time Relative
Proposed 8.497 1.000
L & H 9.317 1.097
Devroye 15.858 1.866
In our view, the data-augmented Gibbs sampler and the exact simulation method complement the existing methods for simulating from the GIG. In the first experiment, we saw that our implementation of the data-augmented Gibbs sampler can be more efficient than the R package boodist for simulating from the GIG. In the second experiment, we saw that our implementation of the exact simulation method can be more efficient than both boodist and GIGrvg when $|p|$ is sufficiently small. Code to replicate these experiments is available at https://github.com/michaeljauch/gig.